OpenStudy (anonymous):

(-8,0) and (0,7) what is the equation of the line (-8,0) and (0,7) what is the equation of the line @Mathematics

OpenStudy (anonymous):

use Y2-Y1 over X2- X1 7-0/0-(-8)=7/8 m=7/8 y=7/8x+b choose any ordered pair and substitute to get your y

OpenStudy (anonymous):

y=7/8x+7

OpenStudy (anonymous):

this is the equation of the line